OpenStudy (whovianchick):

Pre-Calculus help! Will fan and medal! Use synthetic division to find the zeros of the function: f(x)=x^3+x^2+4x+4

OpenStudy (whovianchick):

\[f(x)=x^3+x^2+4x+4\]

zepdrix (zepdrix):

Hey there AvocadoChick! :) You need to apply rational root theorem to find a root before you can apply synthetic division. Our rational root theorem tells us that the factors of 4 will all be possible roots of this polynomial. So we'll have to check each one until we find one that works. We're looking for a root, so we're looking for an x-value that gives us a y-value of zero.\[\large\rm 0=x^3+x^2+4x+4\]

zepdrix (zepdrix):

There are a few factors of 4, 4*1 -4*-1 2*2 -2*-2 Let's start with 1, that's a nice easy number, \(\large\rm x=1:\qquad 1^3+1^2+4(1)+4\ne0\) Hmm that didn't give us zero. Let's try another factor, \(\large\rm x=-1:\qquad (-1)^3+(-1)^2+4(-1)+4=0\) Ooo x=-1 seems to work! If we take this x=-1 and add 1 to each side, we get, \(\large\rm x+1=0\) We found a root! So we can use the root to apply synthetic division.

zepdrix (zepdrix):

Ahh woops, I'm forgetting... We would use x+1 for `polynomial long division`. But simply x=-1 for `synthetic division`. So ignore that last step I applied.

zepdrix (zepdrix):